OpenStudy (anonymous):

limit of ( n/(n-4) )^n as n goes to infinity??

OpenStudy (anonymous):

positive infinitiy

OpenStudy (anonymous):

or wait, nvm

OpenStudy (anonymous):

1

OpenStudy (anonymous):

yea..the calculator says e^4...have no idea how to get there

OpenStudy (anonymous):

\[\lim_{n \rightarrow \infty} [\frac{n}{n-4}]^n\] ?

OpenStudy (anonymous):

yes thats the prob

OpenStudy (anonymous):

are you sure that your answer is e^4?

OpenStudy (anonymous):

because I got 1 too lol

OpenStudy (anonymous):

ok yea thats wat i was thinking too..thank u!

OpenStudy (anonymous):

\[\lim_{n \rightarrow \infty} e^{n \ln \frac{n}{n-4}}\] first you find the limit of the fraction : \[\lim_{n \rightarrow \infty} \frac{n}{n-4} = \lim_{n \rightarrow \infty} \frac{n}{n} = 1\] so now plug this value in the first equation I wrote: \[\lim_{n \rightarrow \infty} e^0 = 1\] np :)

OpenStudy (anonymous):

I get \[e^4\]

OpenStudy (anonymous):

how? .-.

OpenStudy (anonymous):

hmm, I prolly missed out a point that I can't seem to find ._.

OpenStudy (nowhereman):

No, that is not how it works. This way there wouldn't be any constant e. The problem with your approach is, that when the ln goes to zero, at the same time n goes to infinity and so you can't say that their product goes to zero! But you can do the following substitution: \[\lim_{n→∞}{\left(\frac{n}{n-4}\right)^{n}} = \lim_{m→∞}{\left(\frac{m+4}{m}\right)^{m+4}} = \lim_{m→∞}{\left(1+\frac{4}{m}\right)^{m}} \cdot \lim_{m→∞}{\left(\frac{m+4}{m}\right)^{4}} = e^4 \cdot 1

OpenStudy (anonymous):

Set the expression equal to y and take the log of both sides. You'll end up with\[\log y = n \log \frac{n}{n-4}=n \log \frac{1}{1-4/n}=-n \log (1-4/n)=-\frac{\log (1-4/n)}{1/n}\]That's now indeterminate, so you can apply L'Hopital's and take the limit. Then undo the log.

OpenStudy (nowhereman):

\[\lim_{n→∞}{\left(\frac{n}{n-4}\right)^{n}} = \lim_{m→∞}{\left(\frac{m+4}{m}\right)^{m+4}} = \lim_{m→∞}{\left(1+\frac{4}{m}\right)^{m}} \cdot \lim_{m→∞}{\left(\frac{m+4}{m}\right)^{4}}\]\[ = e^4 \cdot 1\]

OpenStudy (anonymous):

when the degree is same on the top and bottom, dont you jus look at the coeffecients of the highest degree which is 1/1= 1

OpenStudy (anonymous):

You should get what nowhereman got.

OpenStudy (anonymous):

/I got.

OpenStudy (nowhereman):

misha: true, but you can only pull the limit inside, if the outer part does not depend on the variable (n here)

OpenStudy (anonymous):

then I was wrong, misha follow both of their ways ^_^

OpenStudy (nowhereman):

You should probably take a look at http://en.wikipedia.org/wiki/Exponential_function or something similar.

OpenStudy (anonymous):

ok thanks for clarifying!

OpenStudy (anonymous):

I gave you a medal sstarica

OpenStudy (anonymous):

lol, thank you :) but I didn't work for it

OpenStudy (anonymous):

for being a good sport :)

OpenStudy (anonymous):

^_^ thank you

OpenStudy (anonymous):

You can have one too, nowhereman...

OpenStudy (anonymous):

lol I gave both of you a medal

OpenStudy (anonymous):

loki

OpenStudy (anonymous):

I prolly am just seeing it, but just to make sure, is there something wrong b/w you and nowhereman?

OpenStudy (anonymous):

Not that I am aware of. Why?

OpenStudy (anonymous):

I've never talked to him.

OpenStudy (anonymous):

mm, lol nvm, you guys just act weird around each other?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

lol, I'm prolly just seeing it, it's nothing ^_^

OpenStudy (anonymous):

he doesn't 'chat'

OpenStudy (anonymous):

noticed :)